Issue $400,000 of bonds. The bonds issue would be developed with a stated rate of 5% and would be a 10 years bong with interest paid semi-annually on June 30 and December 31.

The current market rate for a similar bond is 3%. John would like the journal entry for the bond issue and journal entry for first two interest payments. the company  would use the effective interest rate to amortize any bond discount or premium.
Answer #1

A)Semiannual interest =Par value* stated rate *n/12

                   = 400000 * .05 *6/12

                    = 10000

There are Two semiannual period in a year comprising of 6 months each.

semiannual yield = 3*6/12 = 1.5%

Semiannual period = 10*2 = 20

Issue price = [PVA1.5%,20*Semiannual interest ] +[PVF1.5%,20*Par value]

                = [17.16864*10000] + [.74247*400000]

                = 171686.4+ 296988

                = $ 468,674.4 ( rounded to 468,674)

#Find present value annuity factor from present value annuity table at 1.5% for 20 periods or using financial calculator where PMT=1 ,i=1.5%,n=20

#Find present value factor from present value table at 1.5%,n= 20 or using financial calculator where FV=1,N=20 and i=1.5%

Date Account title Debit credit
1Jan cash 468,674
Premium on bond payable 68674
Bond payable 400000
[Being Bond issued at premium]
30 June Interest expense (468674*1.5%) 7030.11
Premium on bond payable 2969.89
cash 10000
31 Dec Interest expense (465704.11*1.5%) 6985.56
Premium on bond payable 3014.44
cash 10000

#Carrying value of Bond after first semiannual period :Issue price -Premium amortization

                                          = 468674- 2969.89

                                           = 465704.11